# When and How: Learning Identifiable Latent States for Nonstationary Time Series Forecasting

### v1

@@ 5.2.2 Results and Discussion
Quantitative Results. Experiment results on each dataset are shown in Table 2. Please refer to Appendix F.2.2 and F.3 for experiment results on other datasets and sensitive analyis. According to the experiment results, our IDEA model significantly outperforms all other baselines on most forecasting tasks. Specifically, our method outperforms the most competitive baselines by a clear margin of $1.7\%-12\%$ and reduces the forecasting errors substantially on some hard benchmarks, e.g., weather and ILI. Besides outperforming the forecasting model without the nonstationary assumptions like TimesNet and DLinear, our IDEA model also outperforms those methods for nonstationary time series data like RevIN and nonstationary Transformer. However, our method achieves the second-best but still comparable results in the Exchange dataset of the forecasing length of 72, this might be caused by the inaccurate environment estimation for long-term prediction.
It is remarkable that our method achieves a better performance than that of Koopa and SAN, which assume temporal distribution shifts in each time series instance. This is because these methods assume that uniform temporal distribution shifts in each time series instance, which is hard to meet in real-world scenarios, and it is hard for these methods to disentangle the stationary and nonstationary components simultaneously. Meanwhile, our method detects when the temporal distribution shift occurs and further disentangles the stationary and nonstationary states with identification guarantees, hence it can achieve the ideal nonstationary forecasting performance. Please refer to Appendix F.2.2 for experiment results on the M4 dataset.
@@ 5.2.3 Ablation Study
We further devise three model variants. a) IDEA-H: we remove the autoregressive hidden Markov model for environment estimation, and use random environment variables. b) IDEA-E: we remove the nonstationary prior and the corresponding KL divergence term. c) IDEA-S: we remove the stationary prior and the corresponding Kullback-Leibler divergence term. Experiment results on the ILI dataset are shown in Figure 4. We find that 1) the performance of IDEA-H drops without an accurate estimation of the environments, implying that the accurate environment estimation benefits the disentanglement and forecasting performance. 2) Both stationary and non-stationary priors play an important role in non-stationary forecasting, indirectly implying that these priors can benefit the capture of temporal information and the nonstationary forecasting performance.
@@ 6 Conclusion
This paper presents a general data generation process for nonstationary time series forecasting, aligning closely with real-world scenarios. Based on this data generation process, we establish the identifiability of latent stationary latent variables, and nonstationary latent variables via sufficient observation. This contribution stands as a groundbreaking and theoretically sound solution for nonstationary time-series forecasting. Compared with the existing methods, our IDEA model breaks the strong assumption of uniform temporal distribution shifts, making it possible to apply our method to real-world data. Experiment results on several mainstream benchmark datasets further evaluate the effectiveness of our method. In summary, this paper represents a significant stride forward in the field of nonstationary time series forecasting and causal representation learning.
@@ 7 Impact Statement
The proposed IDEA model can detect when the temporal distribution shifts occur and disentangle the stationary and nonstationary latent variables. Therefore, our IDEA could be applied to a wide range of applications including time series forecasting, imputation, and classification. Specifically, the disentangled stationary and nonstationary latent variables would create a model that is more transparent, thereby aiding in the reduction of bias and the promotion of fairness of the existing time series forecasting models.
@@ Appendix A Related Works
Nonstationary Time Series Forecasting. Time series forecasting is a conventional task in the field of machine learning with lots of successful cases, e.g, autoregressive model (Hyndman & Athanasopoulos, 2018) and ARMA (Box & Pierce, 1970). Previously, deep neural networks also have made great contributions to time series forecasting, e.g., RNN-based models (Hochreiter & Schmidhuber, 1997; Lai et al., 2018; Salinas et al., 2020), CNN-based models (Bai et al., 2018; Wang et al., 2022; Wu et al., 2022), and the methods based on state-space model (Gu et al., 2022, 2020, 2021b, 2021a; Smith et al., 2022). Recently, transformer-based methods (Zhou et al., 2021; Wu et al., 2021; Liu et al., 2023a; Nie et al., 2022) have boosted the development of time series forecasting. However, these methods are devised for stationary time series, so nonstationary forecasting is receiving more and more attention. One straightforward solution to this challenge is to discard the nonstationarity via preprocessing methods like stationarization (Virili & Freisleben, 2000) and differencing (Salles et al., 2019), but they might destroy the temporal dependency. Recent studies have used two different assumptions to further solve this problem. By assuming that the temporal distribution shift occurs among datasets and each sequence instance is stationary (Cai et al., 2021; Eldele et al., 2023), some methods consider normalization-based methods. Kim et.al (Kim et al., 2021) propose the reversible instance normalization to remove and restore the statistical information of a time-series instance. Liu et.al (Liu et al., 2022) propose the nonstationary Transformer, which includes the destationary attention mechanism to recover the intrinsic non-stationary information into temporal dependencies. By assuming that the temporal distribution shift uniformly occurs in each sequence instance and so each equal-size segmentation is stationary, other methods propose to disentangle the stationary and nonstationary components. Surana et.al (Surana, 2020) and Liu et.al (Liu et al., 2023b) employ the Koopman theory (Korda & Mezić, 2018), which transform the nonlinear system into several linear operators, to decompose the stationary and nonstationary factors. Liu et.al (Liu et al., 2023c) use adaptive normalization and denormlization on non-overlap equally-sized slices. However, since the temporal distribution shift may occur any time, the aforementioned two assumptions are unreasonable. To solve this problem with milder assumptions, the proposed IDEA first identifies when the distribution shift occurs and then identifies the latent states to learn how they change over time with the help of Markov assumption of latent environment and sufficient observation assumption.
@@
Identifiability of Latent Variables. Identifiability of latent variables (Kong et al., 2023b; Yan et al., 2023; Kong et al., 2023a) plays a significant role in the explanation and generalization of deep generative models, guaranteeing that causal representation learning can capture the underlying factors and describe the latent generation process (Kumar et al., 2017; Locatello et al., 2019a, b; Schölkopf et al., 2021; Träuble et al., 2021; Zheng et al., 2022). Several researchers employ independent component analysis (ICA) to learn causal representation with identifiability (Comon, 1994; Hyvärinen, 2013; Lee & Lee, 1998; Zhang & Chan, 2007) by assuming a linear generation process. To extend it to the nonlinear scenario, different extra assumptions about auxiliary variables or generation processes are adopted to guarantee the identifiability of latent variables (Zheng et al., 2022; Hyvärinen & Pajunen, 1999; Hyvärinen et al., 2023; Khemakhem et al., 2020b; Li et al., 2023c). Previously, Aapo et.al established the identification results of nonlinear ICA by introducing auxiliary variables e.g., domain indexes, time indexes, and class labels(Khemakhem et al., 2020a; Hyvarinen & Morioka, 2016, 2017; Hyvarinen et al., 2019). However, these methods usually assume that the latent variables are conditionally independent and follow the exponential families distributions. Recently, Zhang et.al release the exponential family restriction (Kong et al., 2022; Xie et al., 2022) and propose the component-wise identification results for nonlinear ICA with a certain number of auxiliary variables. They further propose the subspace Identification (Li et al., 2023a) for multi-source domain adaptation, which requires fewer auxiliary variables. In the field of sequential data modeling, Yao et.al (Yao et al., 2021, 2022) recover time-delay latent dynamics and identify their relations from sequential data under the stationary environment and different distribution shifts. And Lippe et.al propose the (i)-CITRIS (Lippe et al., 2022b, a), which use intervention target information for identifiability of scalar and multidimensional latent causal factors. Moreover, Hälvä et.al (Hälvä & Hyvarinen, 2020) and Song et.al (Song et al., 2023) utilize the Markov assumption to provide identification guarantee of time series data without extra auxiliary variables. Although Yao et.al (Yao et al., 2022) partitioned the latent space into stationary and nonstationary parts, they require extra environment variables. Furthermore, although Hälvä et.al (Hälvä & Hyvarinen, 2020) and Song et.al (Song et al., 2023) provide identifiability results without extra environment variables, they can hardly disentangle the stationary and nonstationary, respectively. In this paper, we provide a causal generation process for nonstationary time series data, where the observations are generated by the environment-irrelated stationary latent variables and environment-related nonstationary latent variables. To show the identifiability of these latent variables, we employ the Markov assumption of latent environments and the extension of Kruskal’s theorem (Allman et al., 2009; Kruskal, 1977, 1976) to identify the latent discrete environment variables. Moreover, to disentangle the stationary and nonstationary latent states, we employ the sufficient observation assumption, which requires at least 2 consecutive observations for each latent environment and is reasonable in nonstationary time series, so the symmetry of the stationary and nonstationary latent variables is broken, making it possible to disentangle these latent variables. Summarization and differences of the recent literature related to our work from six different perspectives are shown in Table 3.
Table 3: Summarization and different of the existing identification based on nonlinear ICA. A green check denotes that a method has an attribute, whereas a red cross denotes the opposite.
Theory
Time-varying
Relation
Causally-related
Process
Partitioned
Subspace
Unknown
Environments
Subspace
Identification
TCL (Hyvarinen & Morioka, 2016)
PCL (Hyvarinen & Morioka, 2017)
HM-NLICA (Hälvä & Hyvarinen, 2020)
i-VAE (Khemakhem et al., 2020a)
LEAP (Yao et al., 2021)
TDRL (Yao et al., 2022)
SIG (Li et al., 2023a)
NCTRL (Song et al., 2023)
IDEA
@@ Appendix C Prior Likelihood Derivation
  • We first consider the prior of $\ln p({\mathbf{z}}_{1:t}^{s})$. We start with an illustrative example of stationary latent causal processes with two time-delay latent variables, i.e. ${\mathbf{z}}^{s}_{t}=[z^{s}_{t,1},z^{s}_{t,2}]$ with maximum time lag $L=1$, i.e., $z_{t,i}^{s}=f_{i}({\mathbf{z}}_{t-1}^{s},\epsilon_{t,i}^{s})$ with mutually independent noises. Then we write this latent process as a transformation map $\mathbf{f}$ (note that we overload the notation $f$ for transition functions and for the transformation map):
    $\small\begin{gathered}\begin{bmatrix}\begin{array}[]{c}z_{t-1,1}^{s}\\ z_{t-1,2}^{s}\\ z_{t,1}^{s}\\ z_{t,2}^{s}\end{array}\end{bmatrix}=\mathbf{f}\left(\begin{bmatrix}\begin{array}[]{c}z_{t-1,1}^{s}\\ z_{t-1,2}^{s}\\ \epsilon_{t,1}^{s}\\ \epsilon_{t,2}^{s}\end{array}\end{bmatrix}\right).\end{gathered}$
@@
    $\small p(z_{t-1,1}^{s},z_{t-1,2}^{s},z_{t,1}^{s},z_{t,2}^{s})=\frac{p(z_{t-1,1}^{s},z_{t-1,2}^{s},\epsilon_{t,1}^{s},\epsilon_{t,2}^{s})}{|\text{det }\mathbf{J}_{\mathbf{f}}|},$ (17)
@@
    $\small\begin{gathered}\mathbf{J}_{\mathbf{f}}=\begin{bmatrix}\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ \frac{\partial z_{t,1}^{s}}{\partial z_{t-1,1}^{s}}&\frac{\partial z_{t,1}^{s}}{\partial z_{t-1,2}^{s}}&\frac{\partial z_{t,1}^{s}}{\partial\epsilon_{t,1}^{s}}&0\\ \frac{\partial z_{t,2}^{s}}{\partial z_{t-1,1}^{s}}&\frac{\partial z_{t,2}^{s}}{\partial z_{t-1,2}^{s}}&0&\frac{\partial z_{t,2}^{s}}{\partial\epsilon_{t,2}^{s}}\end{array}\end{bmatrix}.\end{gathered}$
    Given that this Jacobian is triangular, we can efficiently compute its determinant as $\prod_{i}\frac{\partial z_{t,i}^{s}}{\epsilon_{t,i}^{s}}$. Furthermore, because the noise terms are mutually independent, and hence $\epsilon_{t,i}^{s}\perp\epsilon_{t,j}^{s}$ for $j\neq i$ and $\epsilon_{t}^{s}\perp{\mathbf{z}}_{t-1}^{s}$, so we can with the RHS of Equation (17) as follows
    $\small\begin{split}p(z_{t-1,1}^{s},z_{t-1,2}^{s},z_{t,1}^{s},z_{t,2}^{s})=p(z_{t-1,1}^{s},z_{t-1,2}^{s})\times\frac{p(\epsilon_{t,1}^{s},\epsilon_{t,2}^{s})}{|\mathbf{J}_{\mathbf{f}}|}=p(z_{t-1,1}^{s},z_{t-1,2}^{s})\times\frac{\prod_{i}p(\epsilon_{t,i}^{s})}{|\mathbf{J}_{\mathbf{f}}|}.\end{split}$ (18)
@@
    Since the noise $\hat{\epsilon}_{t}^{e}$ is independent of $\hat{{\mathbf{e}}}_{t}$ we have
    $\ln p(\hat{{\mathbf{z}}}_{t}^{e}|\hat{{\mathbf{e}}}_{t})=\ln p(\hat{\epsilon}^{e}_{t})+\sum_{i=1}^{n_{e}}\ln|\frac{\partial r_{i}^{e}}{\partial\hat{z}^{e}_{t,i}}|.$ (23)

### v4

@@ 5.2.2 Results and Discussion
[…]
It is remarkable that our method achieves a better performance than that of FOIL and SOILD, which assume temporal distribution shifts in each time series instance. This is because these methods assume that uniform temporal distribution shifts in each time series instance, which is hard to meet in real-world scenarios, and it is hard for these methods to disentangle the stationary and nonstationary components simultaneously. Meanwhile, our method detects when the temporal distribution shift occurs and further disentangles the stationary and nonstationary states with identification guarantees, hence it can achieve the ideal nonstationary forecasting performance. Please refer to Appendix F.2.2 for experiment results on the M4 dataset.
@@ 5.2.3 Ablation Study
We further devise three model variants. a) UDA-H: We remove the autoregressive hidden Markov model for environment estimation, and use random environment variables. b) UDA-E: We remove the nonstationary prior and the corresponding KL divergence term. c) UDA-S: We remove the stationary prior and the corresponding Kullback-Leibler divergence term. d) UDA-sh: We use a shared decoder for forecasting and reconstruction. Experiment results on the ILI dataset are shown in Figure 5 in Appendix F.4. We find that 1) the performance of UDA-H drops without an accurate estimation of the environments, implying that the accurate environment estimation benefits the disentanglement and forecasting performance. 2) Both stationary and non-stationary priors play an important role in forecasting, implying that these priors can capture temporal information.
@@ 6 Conclusion
This paper introduces the UDA for nonstationary time series forecasting that addresses the challenge of adapting to temporal distribution shifts without relying on the assumption of uniformity. By detecting when distribution shifts occur and disentangling stationary and nonstationary latent variables, the UDA model enables dynamic adaptation to evolving environments. We leverage a Hidden Markov model to identify latent environments and use the variability of historical data to effectively separate stationary and nonstationary components. Through the integration of variational autoencoders and modular prior networks, our model facilitates automatic adaptation to nonstationary changes, significantly enhancing forecasting performance. Experimental results on multiple datasets demonstrate the practical effectiveness and superiority of our approach, marking a key advancement in the field of nonstationary time series forecasting.
@@ 7 Impact Statement
The proposed UDA model can detect when the temporal distribution shifts occur and disentangle the stationary and nonstationary latent variables. Therefore, our UDA could be applied to a wide range of applications including time series forecasting, imputation, and classification. Specifically, the disentangled stationary and nonstationary latent variables would create a model that is more transparent, thereby aiding in the reduction of bias and the promotion of fairness of the existing time series forecasting models.
@@ Appendix A Related Works
Nonstationary Time Series Forecasting. Time series forecasting is a conventional task in the field of machine learning with lots of successful cases, e.g, autoregressive model (Hyndman and Athanasopoulos,, 2018) and ARMA (Box and Pierce,, 1970). Previously, deep neural networks also have made great contributions to time series forecasting, e.g., RNN-based models (Hochreiter and Schmidhuber,, 1997; Lai et al.,, 2018; Salinas et al.,, 2020), CNN-based models (Bai et al.,, 2018; Wang et al.,, 2022; Wu et al.,, 2022), and the methods based on state-space model (Gu et al.,, 2022, 2020; Gu et al., 2021b, ; Gu et al., 2021a, ; Smith et al.,, 2022). Recently, transformer-based methods (Zhou et al.,, 2021; Wu et al.,, 2021; Liu et al., 2023a, ; Nie et al.,, 2022) have boosted the development of time series forecasting.
@@
However, these methods are devised for stationary time series, so nonstationary forecasting is receiving more and more attention (Hu et al.,, 2024). One straightforward solution to this challenge is to discard the nonstationarity via preprocessing methods like stationarization (Virili and Freisleben,, 2000) and differencing (Salles et al.,, 2019), but they might destroy the temporal dependency. Recent studies have used two different assumptions to further solve this problem. By assuming that the temporal distribution shift occurs among datasets and each sequence instance is stationary (Cai et al.,, 2021; Eldele et al.,, 2023), some methods consider normalization-based methods. Kim et al. (Kim et al.,, 2021) propose the reversible instance normalization to remove and restore the statistical information of a time-series instance. Liu et al. (Liu et al.,, 2022) propose the nonstationary Transformer, which includes the destationary attention mechanism to recover the intrinsic non-stationary information into temporal dependencies.
By assuming that the temporal distribution shift uniformly occurs in each sequence instance and so each equal-size segmentation is stationary, other methods propose to disentangle the stationary and nonstationary components. Surana et al. (Surana,, 2020) and Liu et al. (Liu et al., 2023b, ) employ the Koopman theory (Korda and Mezić,, 2018), which transform the nonlinear system into several linear operators, to decompose the stationary and nonstationary factors. Liu et al. (Liu et al., 2023c, ) use adaptive normalization and denormlization on non-overlap equally-sized slices. However, since the temporal distribution shift may occur any time, the aforementioned two assumptions are unreasonable. To solve this problem with milder assumptions, the proposed UDA first identifies when the distribution shift occurs and then identifies the latent states to learn how they change over time with the help of Markov assumption of latent environment and sufficient observation assumption.
@@
Identifiability of Latent Variables. Identifiability of latent variables (Kong et al., 2023b, ; Yan et al.,, 2023; Kong et al., 2023a, ) plays a significant role in the explanation and generalization of deep generative models, guaranteeing that causal representation learning can capture the underlying factors and describe the latent generation process (Kumar et al.,, 2017; Locatello et al., 2019a, ; Locatello et al., 2019b, ; Schölkopf et al.,, 2021; Träuble et al.,, 2021; Zheng et al.,, 2022). Several researchers employ independent component analysis (ICA) to learn causal representation with identifiability (Comon,, 1994; Hyvärinen,, 2013; Lee and Lee,, 1998; Zhang and Chan,, 2007) by assuming a linear generation process. To extend it to the nonlinear scenario, different extra assumptions about auxiliary variables or generation processes are adopted to guarantee the identifiability of latent variables (Zheng et al.,, 2022; Hyvärinen and Pajunen,, 1999; Hyvärinen et al.,, 2023; Khemakhem et al., 2020b, ; Li et al., 2023c, ). Previously, Aapo et al. established the identification results of nonlinear ICA by introducing auxiliary variables e.g., domain indexes, time indexes, and class labels(Khemakhem et al., 2020a, ; Hyvarinen and Morioka,, 2016, 2017; Hyvarinen et al.,, 2019).
However, these methods usually assume that the latent variables are conditionally independent and follow the exponential families distributions. Recently, Zhang et al. release the exponential family restriction (Kong et al.,, 2022; Xie et al.,, 2022) and propose the component-wise identification results for nonlinear ICA with a certain number of auxiliary variables. They further propose the subspace Identification (Li et al., 2023a, ) for multi-source domain adaptation, which requires fewer auxiliary variables. In the field of sequential data modeling, Yao et al. (Yao et al.,, 2021, 2022) recover time-delay latent dynamics and identify their relations from sequential data under the stationary environment and different distribution shifts. And Lippe et al. propose the (i)-CITRIS (Lippe et al., 2022b, ; Lippe et al., 2022a, ), which use intervention target information for identifiability of scalar and multidimensional latent causal factors. Moreover, Hälvä et al. (Hälvä and Hyvarinen,, 2020) and Song et al. (Song et al.,, 2023) utilize the Markov assumption to provide identification guarantee of time series data without extra auxiliary variables. Although Yao et al. (Yao et al.,, 2022) partitioned the latent space into stationary and nonstationary parts, they require extra environment variables. Furthermore, although Hälvä et al. (Hälvä and Hyvarinen,, 2020) and Song et al. (Song et al.,, 2023, 2024) provide identifiability results without extra environment variables, they can hardly disentangle the stationary and nonstationary, respectively.
@@ Appendix C Prior Likelihood Derivation
  • We first consider the prior of $\ln p({\mathbf{z}}_{1:t}^{s})$. We start with an illustrative example of stationary latent causal processes with two time-delay latent variables, i.e. ${\mathbf{z}}^{s}_{t}=[z^{s}_{t,1},z^{s}_{t,2}]$ with maximum time lag $L=1$, i.e., $z_{t,i}^{s}=f_{i}({\mathbf{z}}_{t-1}^{s},\varepsilon_{t,i}^{s})$ with mutually independent noises. Then we write this latent process as a transformation map $\mathbf{f}$ (note that we overload the notation $f$ for transition functions and for the transformation map):
    $\small\begin{gathered}\begin{bmatrix}\begin{array}[]{c}z_{t-1,1}^{s}\\ z_{t-1,2}^{s}\\ z_{t,1}^{s}\\ z_{t,2}^{s}\end{array}\end{bmatrix}=\mathbf{f}\left(\begin{bmatrix}\begin{array}[]{c}z_{t-1,1}^{s}\\ z_{t-1,2}^{s}\\ \varepsilon_{t,1}^{s}\\ \varepsilon_{t,2}^{s}\end{array}\end{bmatrix}\right).\end{gathered}$
@@
    $\small p(z_{t-1,1}^{s},z_{t-1,2}^{s},z_{t,1}^{s},z_{t,2}^{s})=\frac{p(z_{t-1,1}^{s},z_{t-1,2}^{s},\varepsilon_{t,1}^{s},\varepsilon_{t,2}^{s})}{|\text{det }\mathbf{J}_{\mathbf{f}}|},$ (17)
@@
    $\small\begin{gathered}\mathbf{J}_{\mathbf{f}}=\begin{bmatrix}\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ \frac{\partial z_{t,1}^{s}}{\partial z_{t-1,1}^{s}}&\frac{\partial z_{t,1}^{s}}{\partial z_{t-1,2}^{s}}&\frac{\partial z_{t,1}^{s}}{\partial\varepsilon_{t,1}^{s}}&0\\ \frac{\partial z_{t,2}^{s}}{\partial z_{t-1,1}^{s}}&\frac{\partial z_{t,2}^{s}}{\partial z_{t-1,2}^{s}}&0&\frac{\partial z_{t,2}^{s}}{\partial\varepsilon_{t,2}^{s}}\end{array}\end{bmatrix}.\end{gathered}$
    Given that this Jacobian is triangular, we can efficiently compute its determinant as $\prod_{i}\frac{\partial z_{t,i}^{s}}{\varepsilon_{t,i}^{s}}$. Furthermore, because the noise terms are mutually independent, and hence $\varepsilon_{t,i}^{s}\perp\varepsilon_{t,j}^{s}$ for $j\neq i$ and $\varepsilon_{t}^{s}\perp{\mathbf{z}}_{t-1}^{s}$, so we can with the RHS of Equation (17) as follows
    $\small\begin{split}p(z_{t-1,1}^{s},z_{t-1,2}^{s},z_{t,1}^{s},z_{t,2}^{s})=p(z_{t-1,1}^{s},z_{t-1,2}^{s})\times\frac{p(\varepsilon_{t,1}^{s},\varepsilon_{t,2}^{s})}{|\mathbf{J}_{\mathbf{f}}|}=p(z_{t-1,1}^{s},z_{t-1,2}^{s})\times\frac{\prod_{i}p(\varepsilon_{t,i}^{s})}{|\mathbf{J}_{\mathbf{f}}|}.\end{split}$ (18)
@@
    Since the noise $\hat{\varepsilon}_{t}^{e}$ is independent of $\hat{{\mathbf{e}}}_{t}$ we have
    $\ln p(\hat{{\mathbf{z}}}_{t}^{e}|\hat{{\mathbf{e}}}_{t})=\ln p(\hat{\varepsilon}^{e}_{t})+\sum_{i=1}^{n_{e}}\ln|\frac{\partial r_{i}^{e}}{\partial\hat{z}^{e}_{t,i}}|.$ (23)